OpenStudy (anonymous):

Anyone familiar with electrostatics? Need some help. Question and my attempt coming up.

OpenStudy (anonymous):

A uniform line charge \[\rho _{l}=\rho _{o}\] extends from -a to +a along the x-axis. Find the electric field at an arbitrary point P. My attempt: Electric Field eqn. \[E=\frac{ q<r-r \prime> }{ 4\pi \epsilon |r-r'|^{3} }\] r(vector) = \[x _{xhat}+y _{yhat}\] r'(vector)= x' xhat <---- I'm not sure about this.

OpenStudy (anonymous):

You would be much better off calculating the electric potential first.

OpenStudy (anonymous):

Though if the point P is truly arbitrary, the answer is going to be rather complicated. If it's restricted to the y-axis then it's a different story...

OpenStudy (anonymous):

It's arbitrary. I got an insane answer. After that, I have to calculate the potential at an arbitrary point with respect to inf.

OpenStudy (anonymous):

I forgot how to find the potential though.

OpenStudy (anonymous):

\[ \Phi(\vec r) = \int \frac{\rho}{4\pi \epsilon_0 |r-r'| }dr' \]

OpenStudy (anonymous):

If the point was arbitrary, why do I assume that the x-components cancel? If the point lied on the y-axis it would.

OpenStudy (anonymous):

I'll try that Jemurray3.

OpenStudy (anonymous):

Just so I'm clear r = <x,y> r' = <x',0> right?

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

Alright, and I'm not sure if I can do this, but. dr' = d(x' xhat) = xhat dx'. The potential would only point in the x direction, correct?

OpenStudy (anonymous):

The potential doesn't point in any direction, it's a scalar.

OpenStudy (anonymous):

Oh yeah forgot about that. So...what would dr' equal?

OpenStudy (anonymous):

The integral would look like this \[ \Phi = \frac{\rho}{4\pi \epsilon_0} \int_{-a}^a \frac{dx'}{\sqrt{r^2 + x'^2 -2x\cdot x'}}\]

OpenStudy (anonymous):

I think I see what you're saying. x^2+y^2 = r^2. dr' = dx'?

OpenStudy (anonymous):

Now, the key to evaluating this integral will be noting that the bottom is \[ \sqrt{x^2 + y^2 + x'^2 - 2x\cdot x'} = \sqrt{(x-x')^2 + y^2} \] Then, make the substitution u = x-x' , du = -dx' to yield \[ \Phi(\vec{r}) = \frac{\rho}{4\pi \epsilon_0 } \int_{x-a}^{x+a} \frac{du}{\sqrt{u^2+y^2}} \]

OpenStudy (anonymous):

no, r is just the position vector <x,y>, it has nothing to do with polar coordinates or circles.

OpenStudy (anonymous):

You have to be careful with your minus signs. Anyway, moving on from there, you can evaluate this integral via trigonometric substitution. Let \[ u = y\cdot \tan(z), du = y\cdot\sec^2(z) dz\] This yields for the integral \[\int \frac{y\sec^2(z) dz}{y\cdot \sec (z) } = \int \sec(z) dz = \ln|\sec(z) + \tan(z) | \]

OpenStudy (anonymous):

now recall that sec(z) = sqrt( 1 + tan^2(z)) and that tan(z) = u/y. Then we have that the antiderivative becomes \[ \ln|\sqrt{1+u^2/y^2} + u/y| = \ln |(\sqrt{u^2+y^2}+ u)/y| = \ln|\sqrt{u^2+y^2}+u| - \ln|y| \]

OpenStudy (anonymous):

the end of that should be - ln(y). Recall that u goes from (x-a) to (x+a) so plugging those limits in we find that the integral is equal to \[ \ln \left | \frac{\sqrt{(x+a)^2+y^2}+(x+a)}{\sqrt{(x-a)^2+y^2}+(x-a) }\right| \]

OpenStudy (anonymous):

Uhm, holy ****. Did you do that by yourself? I just plugged the integral into wolfram and got that answer.

OpenStudy (anonymous):

Oh good. I solved it in my head as I was typing it in, thanks for confirming.

OpenStudy (anonymous):

I have lots of practice.

OpenStudy (anonymous):

Anyway, that's the potential for this horrible mess. You can get the x and y electric field components just by taking the derivatives with respect to x and y respectively, which would be pretty messy, but hey, you gotta do what you gotta do.

OpenStudy (anonymous):

Pretty effin impressive. Alright, so the potential will be that times k(rho) and the electric field will be the ...negative derivative? If I'm remembering correctly?

OpenStudy (anonymous):

that's correct.

OpenStudy (anonymous):

Alright, cool thank you so much.

OpenStudy (anonymous):

No problem. I'm going to keep going with this and take the derivative with respect to y and then show you something mildly cool, which can help you check your answers to these problems.

OpenStudy (anonymous):

I can just plug them into wolfram if you want me to.

OpenStudy (anonymous):

Okay SO \[ \frac{\partial}{\partial y} \sqrt{(x\pm a) + y^2} +(x\pm a) = \frac{y}{\sqrt{(x\pm a)^2 + y^2}}\]

OpenStudy (anonymous):

Yeah, so could I, but it's just to illustrate a point :)

OpenStudy (anonymous):

Alright, lol

OpenStudy (anonymous):

The partial of y takes a lot of steps and I end up with a very long answer when plugging it in.

OpenStudy (anonymous):

You don't HAVE to do it if you don't want. You've helped a lot already.

OpenStudy (anonymous):

splitting the logarithm up and taking the derivative of each part (we can save the minus sign for later) gives \[ \frac{\partial}{\partial y} \log( \sqrt{(x\pm a)^2 + y^2}+(x\pm a)) = \] \[ \frac{1}{\sqrt{(x\pm a)^2 +y^2} + (x\pm a)} \cdot \frac{y}{\sqrt{(x\pm a)^2 + y^2} }\] so putting in the minus sign and subtracting gives \[ \frac{1}{\sqrt{(x- a)^2 +y^2} + (x- a)} \cdot \frac{y}{\sqrt{(x- a)^2 + y^2} }-\] \[\frac{1}{\sqrt{(x+ a)^2 +y^2} + (x+a)} \cdot \frac{y}{\sqrt{(x+a)^2 + y^2} }\] That's really gross. However, what happens when a -> infinity? Taking the limit gives you the answer \[\vec{E} = \frac{k\rho}{y}\hat y = \frac{\rho}{2\pi\epsilon_0 y} \hat y\] which you should recognize as the electric field from an infinite line charge.

OpenStudy (anonymous):

That's interesting, but is that the answer? Am I supposed to assume that a goes to infinity?

OpenStudy (anonymous):

No, it's a cross check. You would obviously want the limit as a->infinity to produce an infinite line charge, and the limit as a-> zero to produce a point charge. When you're doing complicated problems it's a good idea to take quick limits to check things like this.

OpenStudy (anonymous):

Alright, that makes sense. I guess that's what they mean when they say 'sanity check'? Thank you for your help, man. I wish I could give you more medals lol

OpenStudy (anonymous):

that's exactly what they mean. No problem.